OpenStudy (anonymous):

Solve by elimination method. 0.3x-0.2y=4 0.5x+0.5y=-45/17?

OpenStudy (anonymous):

Hi there. Would you like to learn the material or just get an answer?

OpenStudy (anonymous):

get an anwser

OpenStudy (anonymous):

OK. In that case you can just type both equations into wolfram alpha and it will solve them for you. Like this: http://www.wolframalpha.com/input/?i=0.3x-0.2y%3D4+and+0.5x%2B0.5y%3D-45%2F17

OpenStudy (lilg132):

why not just learn it so you wont ask this question again and you will be able to solve it your self?

OpenStudy (anonymous):

i know elimination method am jst havin a hard time with this question

OpenStudy (lilg132):

anyways make the one of the variables the same value by multiplying the two equations cancel that same value variable out and then solve one of the variables. After getting the value to that variable use it in any of the equations to solve the other variable simple! remember to click good answer

OpenStudy (anonymous):

multiply equn. 1 by 5 and equn. 2 by 2 1.5x - y = 20 x + y = -90/17 add these two equations to eliminate y 2.5x = 20 - 90/17 solve this for x then plug this value of x into second equation to find value of y

OpenStudy (lilg132):

0.3x-0.2y=4 (multiply by 5) 0.5x+0.5y=-45/17 or 2.6470588235294117 (multiply by 2) 1.5x - 1y = 20 1x + 1y = 5.294117647058822 add 2.5x = 25.294117647058822 2.5x / 2.5 = 25.294117647058822 / 2.5 x = ? now use this to work out y can you do the rest from here?

OpenStudy (anonymous):

lemmie try

OpenStudy (lilg132):

can you remember to click good answer to the people who try helping you